OpenStudy (anonymous):

Which of the following is a factor of 64x3 + 343?

OpenStudy (anonymous):

Options?

OpenStudy (anonymous):

4x – 7 None of the above 16x2 – 28x + 49 16x2 + 56x + 49

OpenStudy (anonymous):

Which of the following is a factor of 2x3 – 2?

OpenStudy (anonymous):

For the first one, you must use a^3 + b^3 identity.

OpenStudy (mathmate):

Hint: 1. 64x^3+343 = (4x)^3+7^3 2. x^3+y^3=(x+y)(x^2-xy+y^2)

OpenStudy (anonymous):

3rd option

OpenStudy (anonymous):

where are the options for the 2nd question?

OpenStudy (anonymous):

4x – 7 None of the above 16x2 – 28x + 49 16x2 + 56x + 49

OpenStudy (anonymous):

I asked for the question Which of the following is a factor of 2x3 – 2?

OpenStudy (anonymous):

idk sorry

OpenStudy (anonymous):

For ur 1st question answer is option 3. That is 16x2 – 28x + 49. U mean no options were given for that question.

OpenStudy (anonymous):

x2 + x + 1 x + 1 None of the above x2 + 2x + 1 here it is

OpenStudy (anonymous):

Ist option is the right answer.

OpenStudy (anonymous):

thank you alot... do u happen to know this one? Which of the following is a factor of 8x2 + 21x + 10? a.None of the above b.4x + 5 c.x + 2 d.2x + 5

OpenStudy (anonymous):

(x+2)

OpenStudy (anonymous):

Identify the degree of the polynomial 6xy2 – xy + 8 + 12y.

OpenStudy (anonymous):

Which of the following represents the area of a rectangle whose length is x + 1 and whose width is x + 11? a. x2 + 12x + 11 b.x2 + 11 c.x2 + 10x + 11 d.x2 + 11x + 12

OpenStudy (anonymous):

Factor completely 3x3 – 21x2 – 27x a. 3x(x2 – 7x – 9) b.3(x2 – 7x – 9) c.3x(x + 1)(x – 9) d.3(x3 – 21x2 – 27x)

OpenStudy (mathmate):

Hints: 1. The degree of a polynomial corresponds to the term with the highest degree, including all variables. So in 6xy2 – xy, 6xy^2, the degree is 3, and -xy has a degree of 2. Expand (x+1)(x+11) (using FOIL if necessary to find the answer. 3. 3x3 – 21x2 – 27x note the common factor of 3x, which makes 3x3 – 21x2 – 27x=3x(x^2-7x-9). Don't think there are more rational factors.